# Time-domain study of coupled collective excitations in quantum materials

### v1

@@ III Time-domain probes of collective excitations
To explore collective excitations in the time domain and to study their couplings, a variety of advanced ultrafast tools have been developed, each with distinct advantages. Here, we briefly introduce a few selected techniques that serve as the workhorse in the experiments we reviewed in this article. A schematic for each category of techniques is depicted in Fig. 1b–e. For a typical table-top time-resolved probe introduced below, a femtosecond laser pulse is split into two beams, one of which is used to generate a tailored pump pulse to drive the nonequilibrium dynamics while the other is used to produce a probe pulse, which is either a photon or electron pulse. The frequency of the pump pulse can range from sub-terahertz up to ultraviolet and its polarization can be tuned anywhere between linear and circular, depending on the target excitation.
@@ III.2 Ultrafast electron and X-ray scattering
It is worth noting that the absence of coherent phonons does not imply the absence of time-oscillatory features in a diffraction measurement. Without a net displacement of ions, the pump light can couple to the lattice to the second order in the phonon amplitude, where pairs of phonons at are generated to conserve the total momentum, leading to oscillations of the variance in the atomic displacement at twice the frequency of the phonon mode itself 118. This second-order effect has been demonstrated as intensity oscillations in either Bragg peaks 119 or diffuse scattering 120, and is also known as squeezed phonons 121, 122, 120, 123.
Another rapidly developing area that rides on recent progress in the X-ray free-electron laser technology is time-resolved resonant inelastic X-ray scattering (RIXS), which not only reveals phonon dynamics but can also access charge, orbital, and spin excitations as well as high-order, multi-particle correlation functions that describe coupled spin/charge/orbital excitations 124, 38, 125, 39. With ongoing effort on entering the sub-10-meV energy resolution in RIXS 126, this emerging technique provides a unique window into the nonequilibrium states in correlated materials through dynamical signatures across time, momentum, and energy axes.
@@ III.3 Time-resolved photoemission
The advent of momentum microscope in tr-ARPES experiments has also recently made it feasible to investigate micrometer-sized samples that were challenging to access before. For example, tr-ARPES experiments 131, 132, 89 have demonstrated the ability to probe dark excitons and moiré excitons in monolayer and heterobilayer transition metal dichalcogenides, respectively, whereas they were hard to probe by traditional photon-absorption or photon-emission spectroscopy. As illustrated in the bottom panel of Fig. 1d, only the occupied state at the valence band top can be visualized before the optical pump. With resonant photoexcitation at an exciton energy, photoemission signals from excitons can be observed upon the arrival of the pump pulses 131. The exciton photoemission signals feature an energy-momentum distribution centered below the conduction band bottom with a gap of binding energy (green dispersion line), conforming to theoretical expectations 133.
@@ III.4 Time- and space-resolved probes
Time-resolved spectroscopic and diffraction methods introduced thus far can only provide spatially-averaged information inside the probed region, whereas in general ultrafast dynamics may vary in different regions due to different local environments. Therefore, techniques with both spatial and temporal resolutions play an important role in studying inhomogeneous effects and near-field distribution during ultrafast light-matter interaction. Techniques such as ultrafast electron microscopy 134, 135, photoinduced near-field electron microscopy (PINEM) 136, 137, coherent correlated X-ray imaging 138, 139, 140, and lightwave-driven terahertz scanning tunneling microscopy 141, 142 have all shown a great potential in studying heterogeneity and spatially-varying dynamics present in quantum materials.
As an example to illustrate the relation to collective excitations discussed in this article, ultrafast PINEM exhibits exceptional capabilities in studying the evolution of light-induced plasmonic field distribution 143 (see Fig. 1e), and we will return to specific examples in Sec. IV.5. PINEM is based on ultrafast electron microscopy, which typically uses elastically scattered electrons for imaging. In the PINEM mode, inelastically scattered electrons are used instead, where electrons can absorb and emit spatiotemporally overlapped photons and produce discrete peaks with a constant energy interval in the electron energy loss spectrum. Using the energy-filtered electrons for imaging, one can obtain information about dressed photons such as the near-field distribution of surface plasmon-polaritons 144 and propagation of phonon-polaritons 145. By focusing pump and probe pulses at distinct points and spatially scanning the probe electron pulse, the near-field distributions with space, time, and energy resolution can also be obtained 146.
@@
IV Couplings between collective excitations at non-equilibrium
@@ IV.1 Dressing photons with phonons: phonon-polariton
When solids are exposed to external electromagnetic waves, many dressed states or hybridized quasiparticles emerge. Polaritons are those effects that arise from a strong light-matter interaction between photons and dipole-active excitations. Polaritons can be of various types 55, 56 depending on which excitations they hybridize with. For instance, light can hybridize with magnons, phonons, excitons, and plasmons, forming magnon-polaritons, phonon-polaritons, exciton-polaritons, and plasmon-polaritons (Fig. 1). One hallmark feature of these interactions is the appearance of an avoided crossing between the dispersion of two quasiparticles 58, as illustrated in Fig. 2a. The size of the gap is a measure of the interaction strength, where a larger gap suggests a stronger interaction.
@@
Here we focus on recent ultrafast studies of phonon-polaritons, which arise from the interaction between lattice vibration and electromagnetic waves. When light passes through a crystal, it can drive a transverse optical phonon mode if the frequency of the electromagnetic wave resonates with the phonon, which leads to dipole oscillations and re-radiates the electromagnetic wave. The re-radiated waves can further excite the dipolar vibrations, resulting in the propagation of hybridized light and phonon, namely, phonon-polaritons. As the hybridized quasiparticles inherit the attributes from both light and phonons, their frequencies lie in the frequency range between gigahertz and tens of terahertz, showing great potential for applications in THz spectroscopy, THz imaging, and high-bandwidth signal processing 149.
@@
Many early ultrafast experiments of phonon-polaritons focused on the study of ferroelectrics 150, in which photoexcited coherent phonon-polaritons were used as a probe for low-frequency anharmonic phonons that play a crucial role in driving structural and ferroelectric phase transition near the critical point. Raman spectroscopy was first used to measure the frequency of phonon-polaritons 151 but lacked phase information of their wavepacket. With the development of femtosecond laser, ultrafast optical spectroscopy with spatially distinct pump and probe techniques has enabled us to directly measure the phonon-polaritons waveform in the time domain. In one experimental scheme, broadband polar phonon-polariton modes are produced by illuminating a femtosecond laser pulse onto the surface of a ferroelectric material, where the generation mechanism is based on impulsive stimulated Raman scattering (ISRS) through mixing different frequency components contained within the pulse bandwidth 152, 153. However, this method does not generate phonon-polariton mode with a well-defined frequency. To solve this problem, a four-wave mixing method has been developed 154, 147, 155, 156. In this configuration, two identical, spatiotemporally overlapped femtosecond pulses are used to launch the phonon-polaritons in a nonlinear crystal. The two pump pulses form a transient grating due to interference, and they selectively excite phonon-polaritons with a well-defined scattering wavevector that can be tuned by varying the incidence angle between the pump pulses. As the photoexcited phonon-polariton field modulates the index of refraction that in turn affects the scattering of a separate probe pulse, the periodic probe signal can reveal the presence of phonon-polaritons in a phase-resolved manner. Figure 2b shows the measured dispersion relation of phonon-polaritons in LiNbO3 using this four-wave mixing scheme 147, where five branches and four avoided crossings were observed, showing strong couplings between the photons and phonons in this material.
@@
Experiments using the four-wave mixing methodology have some limitations: ISRS can only excite Raman-active modes 153, while to the lowest order only IR-active phonon modes can couple with electromagnetic waves to form phonon-polaritons. Therefore, the scheme is only feasible for noncentrosymmetric materials where IR phonons are also Raman-active, such as in LiNbO3 and LiTaO3. In addition, these experiments need careful adjustment of the geometry of the two pump beams in order to excite phonon-polaritons with a selected wavevector. An alternative method that overcomes these limitations is time-domain THz spectroscopy 157, where phonon-polaritons are excited by a single beam of THz radiation instead of two beams of visible or near-infrared light pulses. However, in a typical time-domain THz experiment that measures phonon-polariton dispersion, a reference crystal is required for accurate electro-optic sampling to retrieve the phase information across the spectrum of the THz pulse, hence complicating the data analysis and interpretation.
@@
Recently, Luo and coworkers developed a time-of-flight detection method that overcame these challenges and was applied to probing phonon-polaritons in van der Waals magnetic materials 148, including NiI2 and MnPS3. In this scheme, a terahertz pulse centered around 3 THz was used to excite phonon-polaritons while the SHG signal of a 800-nm laser pulse was used as a probe to capture phonon-polaritons before or after they are reflected from the back surface of the sample. More specifically, when the THz pump pulse arrived at the sample surface, multiple oscillatory responses due to phonons and magnons were observed (Fig. 2c, window A). After the initial decay of the oscillations, there was a sudden amplitude revival at around 6 ps, suggesting a reflection of the propagating phonon-polariton wavepacket from the sample back surface (Fig. 2c, insets). After this revival, the signal frequency was also observed to increase towards a longer time delay (Fig. 2c, windows B and C). A clearer view of this phenomenon can be seen in the wavelet transform spectrum in Fig. 2d, which shows the time-dependent signal frequencies. Around time zero, as defined by the THz pulse incidence, the excitation spectrum and the second harmonic frequencies of the THz pump can be seen, including the photoexcited and phonon modes. After 6 ps, a chirped mode emerges, whose frequency increases from 2 to 4 THz over time and it corresponds to the phonon-polariton wavepacket. Given the known sample thickness, the time-dependent frequency spectrum in Fig. 2d can yield the group velocity at different frequencies, which in turn helps one reconstruct the energy-momentum dispersion of the phonon-polariton. Furthermore, the time-dependent attenuation of the phonon-polariton during its propagation allows one to compute the imaginary part of the dispersion, whose divergence near the phonon resonance conforms to the expectation that light quickly dissipates its energy through the phonon mode.
@@
While the aforementioned experiments can detect the presence and dispersion of phonon-polariton wavepackets in the probed volume, they do not reveal the spatial structure of these hybridized quasiparticles. To image the spatiotemporal propagation of phonon-polaritons, ultrafast optical or electron microscopy was used. Nelson and colleagues first demonstrated direct imaging of optical phonon-polariton wavepackets in LiTaO3 using ultrafast optical microscopy that leveraged four-wave mixing 155, 156. In one of the experiments, an anisotropic spatial profile of the pump pulse was applied (see Fig. 3a), which excited phonon-polaritons that propagated mainly along the horizontal direction of the sample 155. The spot size of the probe pulse was significantly larger than the excitation region, so spatial information of the propagation dynamics was readily acquired. The left panel of Fig. 3b shows a typical snapshot of the wavepacket distribution at 1335 fs after the pump incidence. After averaging over the vertical direction in each image captured at different time delays, the spatiotemporal propagation of phonon-polaritons in LaTiO3 was obtained in the right panel of Fig. 3b. Under this experimental framework, the phonon-polariton response can be coherently controlled via spatiotemporal shaping of the optical pump pulses 149. For a more comprehensive survey on the coherent control of phonon-polaritons and THz polaritonics, we refer readers to ref. 149.
@@
Apart from optical probes, electrons have also been successfully deployed to image phonon-polariton propagation dynamics in van der Waals materials 145, leveraging the spatial and spectral information of PINEM. As shown in Fig. 3c, an optical pump pulse first illuminates the edge of a hexagonal boron nitride (hBN) and launches multi-branch phonon-polariton wavepackets. The probing free electrons interact with the propagating phonon-polaritons, broadening the electron energy loss profile due to inelastic scattering (Fig. 3d). In this experiment by Kurman et al., it should be noted that the energy of each phonon-polariton is smaller than the spectral resolution of the instrument. Hence, instead of showing discrete peaks drawn in the schematic in Fig. 1e, the electron energy loss spectrum profile is broadened. Figure 3e shows the averaged energy-filtered electron counts along the edge direction (blue arrow in the inset of Fig. 3e) as a function of distance from the edge (red arrow in the inset of Fig. 3e), indicating the field profile of phonon-polariton along the propagation direction. By studying the distance-time relation of the traveling wavepackets (Fig. 3f), one can observe both acceleration (purple arrow) and deceleration (red arrow) during the propagation of the photoexcited phonon-polaritons. The observed acceleration and deceleration result from the excitation of different spectral components at different times during the pump-hBN interaction. Interestingly, Kurman et al. also showed that the accelerating and decelerating propagation can be modified by changing the exciting photon wavelength and sample thickness 145.
@@
Note that the two featured examples of spatiotemporal imaging differ in the launching mechanism of the phonon-polaritons. While ultrafast optical microscopy used two near-infrared pulses to form a transient grating to excite phonon-polaritons based on the ISRS mechanism 155, the PINEM experiment directly used a mid-IR pulse to coherently excite them 145. Hence, the former methodology was deployed to noncentrosymmetric materials while the latter was used to directly excite dipole-active IR modes to launch the phonon-polaritons. In both cases, the additional spatial information has deepened our understanding of how phonon-polaritons propagate in materials, which will be especially helpful when one considers spatially heterogeneous samples in a real device setting 149, 158, 159.

### v2

@@ III Time-domain probes of collective excitations
To explore collective excitations and their couplings in the time domain, a variety of advanced ultrafast tools have been developed, each with distinct advantages. Here, we briefly introduce a few selected techniques that serve as the workhorses in the experiments we reviewed in this article. A schematic for each category of techniques is depicted in Fig. 1b–e. For a typical table-top time-resolved probe introduced below, a femtosecond laser pulse is split into two beams, one of which is used to generate a tailored pump pulse to drive the nonequilibrium dynamics while the other is used to produce a probe pulse, which is either a photon or an electron pulse. The frequency of the pump pulse can range from sub-terahertz up to ultraviolet and its polarization can be tuned anywhere between linear and circular, depending on the target excitation.
@@ III.2 Ultrafast electron and X-ray scattering
It is worth noting that the absence of coherent phonons does not imply the absence of time-oscillatory features in a diffraction measurement. Without a net displacement of ions, the pump light can couple to the lattice to the second order in the phonon amplitude, where pairs of phonons at are generated to conserve the total momentum, leading to oscillations of the variance in the atomic displacement at twice the frequency of the phonon mode itself 118. This second-order effect induces intensity oscillations in either Bragg peaks 119 or diffuse scattering 120, and is also known as squeezed phonons 121, 122, 120, 123.
Another rapidly developing area that rides on recent progress in the X-ray free-electron laser technology is time-resolved resonant inelastic X-ray scattering (RIXS), which not only reveals phonon dynamics but can also access charge, orbital, and spin excitations as well as high-order, multi-particle correlation functions that describe coupled spin/charge/orbital excitations 124, 38, 125, 39. With ongoing effort towards entering the sub-10-meV energy resolution in RIXS 126, this emerging technique provides a unique window into the nonequilibrium states in correlated materials through dynamical signatures across time, momentum, and energy axes.
@@ III.3 Time-resolved photoemission
The advent of momentum microscopes in tr-ARPES experiments has also recently made it feasible to investigate micrometer-sized samples that were challenging to access before. For example, tr-ARPES experiments 131, 132, 89 have demonstrated the ability to probe dark excitons and moiré excitons in monolayer and heterobilayer transition metal dichalcogenides, respectively, whereas they were hard to probe by traditional photon-absorption or photon-emission spectroscopy. As illustrated in the bottom panel of Fig. 1d, only the occupied states at the valence band top can be visualized before the optical pump. With resonant photoexcitation at an exciton energy, photoemission signals from excitons can be observed upon the arrival of the pump pulses 131. The exciton photoemission signals feature an energy-momentum distribution centered below the conduction band bottom with a gap of binding energy (green dispersion line), in agreement with theoretical expectations 133, 134, 135, 136, 137.
@@ III.4 Time- and space-resolved probes
Time-resolved spectroscopic and diffraction methods introduced thus far can only provide spatially-averaged information inside the probed region, whereas in general ultrafast dynamics may vary in different regions due to different local environments. Therefore, techniques with both spatial and temporal resolutions play an important role in studying inhomogeneous effects and near-field distribution during ultrafast light-matter interactions. Techniques such as ultrafast electron microscopy 138, 139, photoinduced near-field electron microscopy (PINEM) 140, 141, coherent correlated X-ray imaging 142, 143, 144, and lightwave-driven terahertz scanning tunneling microscopy 145, 146 have all shown great potential in studying heterogeneity and spatially-varying dynamics present in quantum materials.
As an example to illustrate the relation to collective excitations discussed in this article, ultrafast PINEM exhibits exceptional capabilities in studying the evolution of light-induced plasmonic field distribution 147 (see Fig. 1e), and we will return to specific examples in Sec. IV.5. PINEM is based on ultrafast electron microscopy, which typically uses elastically scattered electrons for imaging. In the PINEM mode, inelastically scattered electrons are used instead, where electrons can absorb and emit spatiotemporally overlapped photons and produce discrete peaks with a constant energy interval in the electron energy loss spectrum. Using the energy-filtered electrons for imaging, one can obtain information about dressed photons such as the near-field distribution of surface plasmon-polaritons 148 and propagation of phonon-polaritons 149. By focusing pump and probe pulses at distinct points and spatially scanning the probe electron pulse, the near-field distributions with space, time, and energy resolution can also be obtained 150.
@@
IV Couplings between collective excitations at nonequilibrium
@@ IV.1 Dressing photons with phonons: phonon-polariton
When solids are exposed to external electromagnetic waves, many dressed states or hybridized quasiparticles emerge. Polaritons are those effects that arise from a strong light-matter interaction between photons and dipole-active excitations. Polaritons can be of various types 55, 56 depending on which excitations they hybridize with. For instance, light can hybridize with magnons, phonons, excitons, and plasmons, forming magnon-polaritons, phonon-polaritons, exciton-polaritons, and plasmon-polaritons (Fig. 1). One hallmark feature of these interactions is the appearance of an avoided crossing between the dispersions of two quasiparticles 58, as illustrated in Fig. 2a. The size of the gap is a measure of the interaction strength, where a larger gap suggests a stronger interaction.
@@
Here we focus on recent ultrafast studies of phonon-polaritons, which arise from the interaction between lattice vibrations and electromagnetic waves. When light passes through a crystal, it can drive a transverse optical phonon mode if the frequency of the electromagnetic wave resonates with the phonon, which leads to dipole oscillations and consequent re-radiation of the electromagnetic wave. The re-radiated waves can further excite the dipolar vibrations, resulting in the propagation of hybridized light and phonons, namely, phonon-polaritons. As the hybridized quasiparticles inherit the attributes from both light and phonons, their frequencies lie in the frequency range between gigahertz and tens of terahertz, showing great potential for applications in THz spectroscopy, THz imaging, and high-bandwidth signal processing 153.
@@
Many early ultrafast experiments of phonon-polaritons focused on the study of ferroelectrics 154, in which photoexcited coherent phonon-polaritons were used as a probe for low-frequency anharmonic phonons that play a crucial role in driving structural and ferroelectric phase transition near the critical point. Raman spectroscopy was first used to measure the frequency of phonon-polaritons 155 but lacked phase information of their wavepacket. With the development of femtosecond lasers, ultrafast optical spectroscopy with spatially distinct pump and probe techniques has enabled us to directly measure the phonon-polaritons waveform in the time domain. In one experimental scheme, broadband polar phonon-polariton modes are produced by illuminating the surface of a ferroelectric material with a femtosecond laser pulse. These modes are generated from impulsive stimulated Raman scattering (ISRS) through mixing different frequency components contained within the pulse bandwidth 156, 157. However, this method does not generate phonon-polariton modes with a well-defined frequency. To solve this problem, a four-wave mixing method has been developed 158, 151, 159, 160. In this configuration, two identical, spatiotemporally overlapped femtosecond pulses are used to launch the phonon-polaritons in a nonlinear crystal. The two pump pulses form a transient grating due to interference, and they selectively excite phonon-polaritons with a well-defined scattering wavevector that can be tuned by varying the incidence angle between the pump pulses. As the photoexcited phonon-polariton field modulates the index of refraction that in turn affects the scattering of a separate probe pulse, the periodic probe signal can reveal the presence of phonon-polaritons in a phase-resolved manner. Figure 2b shows the measured dispersion relation of phonon-polaritons in LiNbO3 using this four-wave mixing scheme 151, where five branches and four avoided crossings were observed, showing strong couplings between the photons and phonons in this material.
@@
Experiments using four-wave mixing have some limitations: ISRS can only excite Raman-active modes 157, while to the lowest order only IR-active phonon modes can couple with electromagnetic waves to form phonon-polaritons. Therefore, this scheme is only feasible for noncentrosymmetric materials in which IR phonons are also Raman-active, such as in LiNbO3 and LiTaO3. In addition, these experiments need careful adjustment of the geometry of the two pump beams in order to excite phonon-polaritons with a selected wavevector. An alternative method that overcomes these limitations is time-domain THz spectroscopy 161, where phonon-polaritons are excited by a single beam of THz radiation instead of two beams of visible or near-infrared light pulses. However, in a typical time-domain THz experiment that measures phonon-polariton dispersion, a reference crystal is required for accurate electro-optic sampling to retrieve the phase information across the spectrum of the THz pulse, hence complicating the data analysis and interpretation.
@@
Recently, Luo and coworkers developed a time-of-flight detection method that overcame these challenges and was applied to probing phonon-polaritons in van der Waals magnetic materials 152, including NiI2 and MnPS3. In this scheme, a pulse centered around 3 THz was used to excite phonon-polaritons while the SHG signal of a 800-nm laser pulse was used as a probe to capture phonon-polaritons before or after they were reflected from the back surface of the sample. More specifically, when the THz pump pulse arrived at the sample surface, multiple oscillatory responses due to phonons and magnons were observed (Fig. 2c, window A). After the initial decay of the oscillations, there was a sudden amplitude revival at around 6 ps, suggesting a reflection of the propagating phonon-polariton wavepacket from the sample back surface (Fig. 2c, insets). After this revival, the signal frequency was also observed to increase towards longer time delay (Fig. 2c, windows B and C). A clearer view of this phenomenon can be seen in the wavelet transform spectrum in Fig. 2d, which shows the time-dependent signal frequencies. Around time zero, as defined by the THz pulse incidence, the excitation spectrum and the second harmonic frequencies of the THz pump can be seen, including the photoexcited and phonon modes. After 6 ps, a chirped mode emerges, whose frequency increases from 2 to 4 THz over, which corresponds to the phonon-polariton wavepacket. Given the known sample thickness, the time-dependent frequency spectrum in Fig. 2d can yield the group velocity at different frequencies, which in turn helps reconstruct the energy-momentum dispersion of the phonon-polariton. Furthermore, the time-dependent attenuation of the phonon-polariton during its propagation allows one to compute the imaginary part of the dispersion, whose divergence near the phonon resonance conforms to the expectation that light quickly dissipates its energy through the phonon mode.
@@
While the aforementioned experiments can detect the presence and dispersion of phonon-polariton wavepackets in the probed volume, they do not reveal the spatial structure of these hybridized quasiparticles. To image the spatiotemporal propagation of phonon-polaritons, ultrafast optical or electron microscopy can be used. Nelson and colleagues first demonstrated direct imaging of optical phonon-polariton wavepackets in LiTaO3 using ultrafast optical microscopy that leveraged four-wave mixing 159, 160. In one of the experiments, a pump pulse with an anisotropic spatial profile was applied (see Fig. 3a), which excited phonon-polaritons that propagated mainly along the horizontal direction of the sample 159. The spot size of the probe pulse was significantly larger than the excitation region, so spatial information of the propagation dynamics was readily acquired. The left panel of Fig. 3b shows a representative snapshot of the wavepacket distribution at 1335 fs after the pump incidence. After averaging over the vertical direction in each image captured at different time delays, the spatiotemporal propagation of phonon-polaritons in LaTiO3 was obtained in the right panel of Fig. 3b. Under this experimental framework, the phonon-polariton response can be coherently controlled via spatiotemporal shaping of the optical pump pulses 153. For a more comprehensive survey on the coherent control of phonon-polaritons and THz polaritonics, we refer readers to ref. 153.
@@
Apart from optical probes, electrons have been successfully deployed to image phonon-polariton propagation dynamics in van der Waals materials 149, leveraging the spatial and spectral information of PINEM. As shown in Fig. 3c, an optical pump pulse first illuminates the edge of a hexagonal boron nitride (-BN) and launches multi-branch phonon-polariton wavepackets. The probing free electrons interact with the propagating phonon-polaritons, broadening the electron energy loss profile due to inelastic scattering (Fig. 3d). In this experiment by Kurman et al., it should be noted that the energy of each phonon-polariton is smaller than the spectral resolution of the instrument. Hence, instead of showing discrete peaks drawn in the schematic in Fig. 1e, the electron energy loss spectrum profile is broadened. Figure 3e shows the averaged energy-filtered electron counts along the edge direction (blue arrow in the inset of Fig. 3e) as a function of distance from the edge (red arrow in the inset of Fig. 3e), indicating the field profile of phonon-polaritons along the propagation direction. By studying the distance-time relation of the traveling wavepackets (Fig. 3f), one can observe both acceleration (purple arrow) and deceleration (red arrow) during the propagation of the photoexcited phonon-polaritons. The observed acceleration and deceleration result from the excitation of different spectral components at different times during the pump--BN interaction. Interestingly, Kurman et al. also showed that the accelerating and decelerating propagation can be modified by changing the exciting photon wavelength and sample thickness 149.
@@
Note that the two featured examples of spatiotemporal imaging differ in the launching mechanism of the phonon-polaritons. While ultrafast optical microscopy used two near-infrared pulses to form a transient grating to excite phonon-polaritons based on the ISRS mechanism 159, the PINEM experiment directly used a mid-IR pulse to coherently excite them 149. Hence, the former methodology was deployed to noncentrosymmetric materials while the latter was used to directly excite dipole-active IR modes to launch the phonon-polaritons. In both cases, the additional spatial information has deepened our understanding of how phonon-polaritons propagate in materials, which will be especially helpful when one considers spatially heterogeneous samples in a real device setting 153, 162, 163.